Problem 1a. Verify that the given functions form a fundamental set of solutions of the differential equation.

4y'' + 36y = 0; y1 = cos(3x); y2 = sin(3x)
Answer #1

Arsy Y y +36 9 zo y +99 → (D2+9) y = 0 Ausillary equation » m= 73, m2+9 5o y = S, COS 3x + casins x hence cy, cos3x z and si

answered by: ANURANJAN SARSAM
Answer #2

Arsy Y y +36 9 zo y +99 → (D2+9) y = 0 Ausillary equation » m= 73, m2+9 5o y = S, COS 3x + casins x hence cy, cos3x z and si